calculate mole fraction of solute in 2M Aqueous solution of acetic acid .the density of solution is 1.02gram/ml.
1
Expert's answer
2012-07-03T08:43:47-0400
the solution molarity is
c = n(acetic acid) / V(solution)
the one liter of
solution has mass:
m(solution) = d(solution) * V(solution) = 1.02 g/mL * 1000
mL = 1020 g
the mass of acetic acid is
m(acetic acid) = n(acetic acid) *
M(acetic acid) = 2 mol * 60 g/mol = 120 g
the mass of water is
m(water) =
m(solution) - m(acetic acid) = 1020 - 120 = 900 g
the quantity of water
is
n(water) = m(water) / M(water) = 900 g / 18 g/mol = 50 mol
the mole
fraction of solute (acetic acid) is
x = n(acetic acid) / ( n(acetic acid) +
n(water) ) = 2 mol / 52 mol = 0.0385